OpenStudy (anonymous):

Which of the following statements are true about functions and relations? The vertical line test will not work for piece-wise defined relations. No piece-wise relations can be functions. The vertical line test must only cross the curve one place for each x value. The vertical line test can cross the curve at more than one place for each x value.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

C right?

OpenStudy (anonymous):

@phi can you help me?

OpenStudy (phi):

yes, The vertical line test must only cross the curve one place for each x value.

OpenStudy (anonymous):

@phi did you see all the things that kid was typing?

OpenStudy (phi):

yes, he is/was a troll.

OpenStudy (anonymous):

oh ok then that explains a lot... can you help me with one more problem?

OpenStudy (anonymous):

Which one of the following is a function? 5x – 2y = 10 5x – 2y^2 = 9 5x^2 – 2y^2 = 9

OpenStudy (anonymous):

@e.mccormick

OpenStudy (phi):

any guess? (simplest is the best guess)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

5x – 2y = 10 right? @phi that is a function

OpenStudy (phi):

yes. that is the equation of a line. and lines (except a vertical one) will pass the vertical line test. the others are hyperbolas... the y^2 part cause trouble.

OpenStudy (anonymous):

thanks!!!